”Measurement of B̄^{0} → D^{∗+}ℓ^{−}ν̄ at CLEO”

I present a study of the exclusive semileptonic B-meson decay B̄^{0} → D^{∗+}ℓ^{−}ν̄ carried out by the CLEO collaboration. I fully reconstruct the D^{∗} and charged lepton and determine the angle between the decaying B-meson and the D^{∗} − ℓ combination. I present a preliminary measurement of the branching ratio B(B̄^{0} → D^{∗+}ℓ^{−}ν̄) obtained through a fit to the distribution of this angle. This fit separates our signal decay mode from other B decays involving higher D resonance states or additional hadrons in the final state.I discuss the important systematics errors in this analysis, in particular the slow pion tracking efficiency. Finally, I present the technique which will be used to extract the CKM matrix element V_{c}b with similar methods, and I give the status of that work. This analysis uses a data sample of 3.3 million BB̄ events collected with the CLEO II detector at the Cornell Electron Storage Ring.